Fermi National Laboratory
 

Operation Summary for January 27, 2010 - January 29, 2010

 

Wednesday Jan. 27

The day shift began with the Accelerator Complex conducting maintenance while Linac personnel cleaned the H- Source.

During the day shift, vacuum, LCW, EE Support technicians, and accelerator personnel accessed into the Tevatron (TeV), Main Injector (MI), Booster, and the Antiproton Source (Pbar). Vacuum techs found a MI RF vacuum leak.

During the evening shift, cryo system techs worked on the TeV sector F4 cryo system. Preaccelerator personnel reported that that the H- Source had recovered.

Thursday Jan. 28

The midnight shift began with an access into the TeV and with Preaccelerator work.

During the midnight shift, the Linac H- Source was left to warm up over the night. MI and TeV safety system tests began and ran through the end of shift. Operators prepared Pbar for access.

During the day shift, recovery from the short maintenance period occurred in stages during the day. Operations established beam to Linac by 10:30 AM and to Booster by 11 AM. The MI tunnel access for vacuum work in the RF sector carried on through to the end of shift and inhibited beam to all users. TeV turn on took most of the afternoon. Operations put the TeV through a dry squeeze.

During the evening shift, MI personnel came out of access at 8 PM. Operations had established beam to MiniBooNE by 8:30 PM, to Recycler by 9:30 PM, to NuMI by 10:30 PM, and to the TeV by 11:30 PM. The beam in MI required some tuning before Recycler and NuMI were able to take beam, and the TeV beam was delayed by a lead failure for a quadrupole power supply (C:D0Q2) that was caused by a bad VFC card. Beam to Pbar was inhibited by a beam stop collimator getting stuck in the beam pipe. Operators called in mechanical technicians to investigate; this lasted to the end of shift.

Friday Jan. 29

The midnight shift began with Operations monitoring a Recycler stash (6.8E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Recycler proton studies lasted until 5 AM. TeV orbit checks and proton only studies lasted until 5:30 AM when a D0 quadrupole magnet lead failure led to a quench. The quench recovery and investigation lasted through out the rest of shift. The MiniBooNE and NuMI experiments took protons when available; this beam was inhibited for a few hours by accesses into the Pbar transport tunnel to work on the beam stop collimator, this work lasted until 7 AM. A Linac Klystron RF station PFN developed a LCW leak and inhibited beam for the rest of the shift.

The Plans for Friday and the Weekend

The plans for today are to resume stacking and stashing after the Linac Klystron LCW leak is repaired. The Run Coordinator will allow Recycler personnel to conduct drag rate measurements and for some NuMI work. The Run Coordinator is hoping to have a shot in the TeV before 4 PM. The weekend plans are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory